# Testing Conditional Independence via the Spectral Generalized Covariance Measure: Beyond Euclidean Data

### v2

@@ B.2 Uniformly Asymptotic Properties
Proof of Lemma B.3.
For any $h\in[0,r_{0}]$, the lower bound on $f$ yields
$F\left(q+h\right)-F(q)=\int_{q}^{q+h}f(u)\mathop{}\!\mathrm{d}u\geq m_{0}h,\quad F(q)-F\left(q-h\right)=\int_{q-h}^{q}f(u)\mathop{}\!\mathrm{d}u\geq m_{0}h,$
hence
$F\left(q+h\right)\geq\tau+m_{0}h,\qquad F\left(q-h\right)\leq\tau-m_{0}h.$
If $\|G-F\|_{\infty}\leq\Delta$, then
$G\left(q-\frac{\Delta}{m_{0}}\right)$ $\leq F\left(q-\frac{\Delta}{m_{0}}\right)+\Delta\leq\tau-\Delta+\Delta=\tau,$
$G\left(q+\frac{\Delta}{m_{0}}\right)$ $\geq F\left(q+\frac{\Delta}{m_{0}}\right)-\Delta\geq\tau+\Delta-\Delta=\tau.$
By the definition of the $\tau$-quantile,
@@
which proves $\left|Q_{\tau}(G)-Q_{\tau}(F)\right|\leq\Delta/m_{0}$. ∎
@@ B.3.2 Error from Eigenfunction Truncation
$n_{1}\left|\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}-T_{n_{1}}^{(P_{n},Q_{n})}\right|$ $=\left|n_{1}\left(\|U\|^{2}-\|V\|^{2}\right)\right|$
@@
$\leq n_{1}(\|U\|+\|V\|)\|U-V\|.$
Since orthogonal projectors have an operator norm of at most $1$,
$\left|U\right|\leq|\Sigma_{n_{1}}|,\qquad\left|V\right|\leq|\Sigma_{n_{1}}|,$
@@
$\left|\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}-T_{n_{1}}^{(P_{n},Q_{n})}\right|\leq 2\,n_{1}\,\|\Sigma_{n_{1}}\|\|U-V\|.$ (A.19)
@@
$\Delta_{X,P_{n}}\coloneqq\hat{P}_{X,P_{n}}-P_{X,P_{n}},\quad\Delta_{Y,Q_{n}}\coloneqq\hat{P}_{Y,Q_{n}}-P_{Y,Q_{n}}.$
Then, by bilinearity of the tensor product,
@@
Using submultiplicativity and the tensor-operator identity $\left\|A\otimes B\otimes C\right\|_{\textrm{op}}=\left\|A\right\|_{\textrm{op}}\left\|B\right\|_{\textrm{op}}\left\|C\right\|_{\textrm{op}}$, together with $\left\|P_{X,P_{n}}\right\|_{\textrm{op}}=\left\|P_{Y,Q_{n}}\right\|_{\textrm{op}}=1$, we obtain
$\|U-V\|\leq\left(\left\|\Delta_{X,P_{n}}\right\|_{\textrm{op}}+\left\|\Delta_{Y,Q_{n}}\right\|_{\textrm{op}}+\left\|\Delta_{X,P_{n}}\right\|_{\textrm{op}}\left\|\Delta_{Y,Q_{n}}\right\|_{\textrm{op}}\right)\|\Sigma_{n_{1}}\|.$
@@
$\left|\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}-T_{n_{1}}^{(P_{n},Q_{n})}\right|\leq 2n_{1}\|\Sigma_{n_{1}}\|^{2}\left\{\left\|\Delta_{X,P_{n}}\right\|_{\textrm{op}}+\left\|\Delta_{Y,Q_{n}}\right\|_{\textrm{op}}+\left\|\Delta_{X,P_{n}}\right\|_{\textrm{op}}\left\|\Delta_{Y,Q_{n}}\right\|_{\textrm{op}}\right\}.$ (A.20)
Therefore, since $n_{1}\|\Sigma_{n_{1}}\|^{2}=O_{p}(1)$ under the null, it suffices to show that
$\|\Delta_{X,P_{n}}\|_{\textrm{op}}=o_{p}(1),\qquad\|\Delta_{Y,Q_{n}}\|_{\textrm{op}}=o_{p}(1).$
We detail the $X$-part; the $Y$-part is identical with $(\mu_{j},\gamma^{Y}_{Q_{n}})$ in place of $(\lambda_{j},\gamma^{X}_{P_{n}})$.
@@
$r_{P_{n}}\coloneqq\frac{\lambda_{P_{n}}+\lambda_{P_{n}+1}}{2},\qquad\Gamma_{P_{n}}\coloneqq\{z\in\mathbb{C}:\left|z\right|=r_{P_{n}}\}.$
Then for every $z\in\Gamma_{P_{n}}$,
$\inf_{\mu\in\sigma(C_{XX})}\left|z-\mu\right|\geq\frac{\gamma^{X}_{P_{n}}}{2}\implies\sup_{z\in\Gamma_{P_{n}}}\left\|(zI-C_{XX})^{-1}\right\|_{\textrm{op}}\leq\frac{2}{\gamma^{X}_{P_{n}}}.$
Let $\hat{C}_{XX}$ be the empirical covariance. On the event
$\Omega^{X}_{n}\coloneqq\left\{\,\left\|\hat{C}_{XX}-C_{XX}\right\|_{\textrm{HS}}<\frac{\gamma^{X}_{P_{n}}}{4}\,\right\},$
Weyl’s bound yields $\inf_{\mu\in\sigma(\hat{C}_{XX})}\left|z-\mu\right|\geq\gamma^{X}_{P_{n}}/4$ for all $z\in\Gamma_{P_{n}}$, hence
$\sup_{z\in\Gamma_{P_{n}}}\left\|\left(zI-\hat{C}_{XX}\right)^{-1}\right\|_{\textrm{op}}\leq\frac{4}{\gamma^{X}_{P_{n}}}\qquad\text{on }\Omega^{X}_{n}.$
By the Riesz formula (Kato, 1995, Theorem 6.17),
@@
$P_{X,P_{n}}=\frac{1}{2\pi i}\oint_{\Gamma_{P_{n}}}(zI-C_{XX})^{-1}\mathop{}\!\mathrm{d}z,\qquad\hat{P}_{X,P_{n}}=\frac{1}{2\pi i}\oint_{\Gamma_{P_{n}}}\left(zI-\hat{C}_{XX}\right)^{-1}\mathop{}\!\mathrm{d}z,$
whence
$\hat{P}_{X,P_{n}}-P_{X,P_{n}}=\frac{1}{2\pi i}\oint_{\Gamma_{P_{n}}}\Bigl[\left(zI-\hat{C}_{XX}\right)^{-1}-(zI-C_{XX})^{-1}\Bigr]\mathop{}\!\mathrm{d}z.$
Let
$R\coloneqq(zI-C_{XX})^{-1},\quad\hat{R}\coloneqq(zI-\hat{C}_{XX})^{-1},\quad D\coloneqq\hat{C}_{XX}-C_{XX}.$
Resolvent identity (Conway, 2007, Proposition 3.9. in VII§4) gives
$\hat{R}-R=RD\hat{R}$.
Hence
$\hat{R}-R=RD\hat{R}=RD(R+RD\hat{R})=RDR+RD\hat{R}DR.$
Substituting back,
$(zI-\hat{C}_{XX})^{-1}-(zI-C_{XX})^{-1}$
$\qquad=(zI-C_{XX})^{-1}(\hat{C}_{XX}-C_{XX})(zI-C_{XX})^{-1}$
$\qquad\quad+(zI-C_{XX})^{-1}(\hat{C}_{XX}-C_{XX})(zI-\hat{C}_{XX})^{-1}(\hat{C}_{XX}-C_{XX})(zI-C_{XX})^{-1}.$
to obtain a linear part and a quadratic remainder.
Let $L_{P_{n}}$ denote the length of $\Gamma_{P_{n}}$; since $r_{P_{n}}\leq\lambda_{1}$, we have $L_{P_{n}}=2\pi r_{P_{n}}\leq 2\pi\lambda_{1}\eqqcolon L_{\star}$. On $\Omega^{X}_{n}$,
$\left\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\right\|_{\textrm{op}}\leq\frac{L_{\star}}{2\pi}\left(\frac{2}{\gamma^{X}_{P_{n}}}\right)^{2}\left\|\hat{C}_{XX}-C_{XX}\right\|_{\textrm{op}}\ +\ \frac{L_{\star}}{2\pi}\left(\frac{2}{\gamma^{X}_{P_{n}}}\right)^{2}\left(\frac{4}{\gamma^{X}_{P_{n}}}\right)\left\|\hat{C}_{XX}-C_{XX}\right\|_{\textrm{op}}^{2}.$ (A.21)
Since $\left\|\hat{C}_{XX}-C_{XX}\right\|_{\textrm{op}}\leq\left\|C_{n}-C\right\|=O_{p}\left(n^{-1/2}\right)$ by Lemma B.14, we conclude
$\left\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\right\|_{\textrm{op}}=O_{p}\left(\frac{1}{\gamma^{X}_{P_{n}}\sqrt{n}}+\frac{1}{\left(\gamma^{X}_{P_{n}}\right)^{2}n}\right)=O_{p}\left(\frac{1}{\gamma^{X}_{P_{n}}\sqrt{n}}\right).$
Define $s_{Q_{n}}\coloneqq\frac{\mu_{Q_{n}}+\mu_{Q_{n}+1}}{2}$ and $\Gamma^{\prime}_{Q_{n}}\coloneqq\left\{z\in\mathbb{C}:\left|z\right|=s_{Q_{n}}\right\}$, and the event
@@
$\Omega^{Y}_{n}\coloneqq\left\{\left\|\hat{C}_{YY}-C_{YY}\right\|_{\textrm{HS}}<\frac{\gamma^{Y}_{Q_{n}}}{4}\right\}.$
Repeating the same steps with $(\mu_{j},\gamma^{Y}_{Q_{n}},C_{YY},\hat{C}_{YY},\Gamma^{\prime}_{Q_{n}})$ gives
$\left\|\hat{P}_{Y,Q_{n}}-P_{Y,Q_{n}}\right\|_{\textrm{op}}=O_{p}\left(\frac{1}{\gamma^{Y}_{Q_{n}}\sqrt{n}}\right).$
This completes the required bounds.
For uniform version, it suffices to show that
$\left\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\right\|_{\textrm{op}}=o_{\tilde{\mathcal{P}}_{0}}(1),\quad\left\|\hat{P}_{Y,Q_{n}}-P_{Y,Q_{n}}\right\|_{\textrm{op}}=o_{\tilde{\mathcal{P}}_{0}}(1)$ (A.22)
because if (A.22) holds, then since $\sqrt{n_{1}}\|\Sigma_{n_{1}}\|=O_{\tilde{\mathcal{P}}_{0}}(1)$ by Theorem 2 in Lundborg, Shah and Peters (2022) and Lemma S5 in Lundborg et al. (2024), we have
$\left|\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}-T_{n_{1}}^{(P_{n},Q_{n})}\right|=o_{\tilde{\mathcal{P}}_{0}}(1),$
by (A.20). (A.22) can be proven because under Assumption 4.5
@@
$\sup_{P\in\tilde{\mathcal{P}}}\mathbb{P}_{P}\left(\left\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\right\|_{\textrm{op}}>\epsilon\right)\leq\frac{C}{\epsilon c_{P_{n}}^{X}\sqrt{n_{1}}}\to 0,$
where $C$ is an upper bound independent of $P\in\tilde{\mathcal{P}}_{0}$ due to the boundedness of the kernel. This inequality is obtained from (A.21) and Markov’s inequality. ∎

### v3

@@ B.2 Uniformly Asymptotic Properties
Proof.
Set $q\coloneqq Q_{\tau}(F)$. Since $F$ admits a density $f$ on $q-r_{0},q+r_{0}]$ with $f\geq m_{0}>0$ there, $F$ is continuous and strictly increasing on $q-r_{0},q+r_{0}]$. In particular, $F(q)=\tau$ (otherwise, if $F(q)>\tau$, continuity would yield some $t<q$ with $F(t)\geq\tau$, contradicting the definition of $q$).
Fix $\Delta$ with $\|G-F\|_{\infty}\leq\Delta<m_{0}r_{0}$ and define $h\coloneqq\Delta/m_{0}\in(0,r_{0})$. The lower bound on $f$ yields
$F(q+h)-F(q)=\int_{q}^{q+h}f(u)\,\mathop{}\!\mathrm{d}u\geq m_{0}h=\Delta,\qquad F(q)-F(q-h)=\int_{q-h}^{q}f(u)\,\mathop{}\!\mathrm{d}u\geq m_{0}h=\Delta.$
Since $F(q)=\tau$, we obtain
$F(q+h)\geq\tau+\Delta,\qquad F(q-h)\leq\tau-\Delta.$
Using $\|G-F\|_{\infty}\leq\Delta$, it follows that
$G(q+h)\geq F(q+h)-\Delta\geq\tau,$
and therefore $Q_{\tau}(G)\leq q+h$ by the definition of the $\tau$-quantile.
For the lower bound, let t<qht<q-h. Then $t\leq q-r_{0}$ need not hold, but we may still compare $F(t)$ to F(qh)F(q-h) by splitting the integral:
$F(q-h)-F(t)=\int_{t}^{q-h}f(u)\,\mathop{}\!\mathrm{d}u\geq\int_{q-h}^{q-h}f(u)\,\mathop{}\!\mathrm{d}u+\int_{t}^{q-h}f(u)\,\mathop{}\!\mathrm{d}u.$
Since $f\geq m_{0}$ holds on $q-r_{0},q+r_{0}]$ and $q-h\in[q-r_{0},q+r_{0}]$, we have
$F(q-h)-F(t)\geq\int_{q-h}^{q-h}f(u)\,\mathop{}\!\mathrm{d}u+\int_{t}^{q-h}0\,\mathop{}\!\mathrm{d}u\geq 0,$
and more importantly, for any t<qht<q-h,
$F(q-h)-F(t)\geq\int_{q-h}^{q-h}f(u)\,\mathop{}\!\mathrm{d}u+\int_{t}^{q-h}0\,\mathop{}\!\mathrm{d}u\geq m_{0}\,(q-h-t),$
where the last inequality uses that on the portion of the interval lying inside $q-r_{0},q+r_{0}]$ the integrand is at least $m_{0}$, while outside it the integrand is nonnegative. Hence,
$F(t)\leq F(q-h)-m_{0}(q-h-t)\leq(\tau-\Delta)-m_{0}(q-h-t)<\tau-\Delta.$
Consequently,
$G(t)\leq F(t)+\Delta<\tau.$
Since $G(t)<\tau$ for all t<qht<q-h, we conclude that $Q_{\tau}(G)\geq q-h$.
Combining the two bounds gives
@@
and hence $\left|Q_{\tau}(G)-Q_{\tau}(F)\right|\leq\Delta/m_{0}$. ∎
@@ B.3.2 Error from Eigenfunction Truncation
$n_{1}\left|\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}-T_{n_{1}}^{(P_{n},Q_{n})}\right|$ $=n_{1}\left|\left\|U\right\|^{2}-\left\|V\right\|^{2}\right|$
@@
$\leq n_{1}\left(\left\|U\right\|+\left\|V\right\|\right)\left\|U-V\right\|.$
Since orthogonal projectors have operator norm at most $1$,
$\left\|U\right\|\leq\left\|\Sigma_{n_{1}}\right\|,\qquad\left\|V\right\|\leq\left\|\Sigma_{n_{1}}\right\|,$
@@
$n_{1}\left|\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}-T_{n_{1}}^{(P_{n},Q_{n})}\right|\leq 2\,n_{1}\,\left\|\Sigma_{n_{1}}\right\|\,\left\|U-V\right\|.$ (A.19)
@@
$\Delta_{X,P_{n}}\coloneqq\hat{P}_{X,P_{n}}-P_{X,P_{n}},\qquad\Delta_{Y,Q_{n}}\coloneqq\hat{P}_{Y,Q_{n}}-P_{Y,Q_{n}}.$
By bilinearity of the tensor product,
@@
Using submultiplicativity and $\left\|A\otimes B\otimes C\right\|_{\mathrm{op}}=\left\|A\right\|_{\mathrm{op}}\left\|B\right\|_{\mathrm{op}}\left\|C\right\|_{\mathrm{op}}$ together with $\left\|P_{X,P_{n}}\right\|_{\mathrm{op}}=\left\|P_{Y,Q_{n}}\right\|_{\mathrm{op}}=1$, we obtain
$\left\|U-V\right\|\leq\left(\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}+\left\|\Delta_{Y,Q_{n}}\right\|_{\mathrm{op}}+\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}\left\|\Delta_{Y,Q_{n}}\right\|_{\mathrm{op}}\right)\left\|\Sigma_{n_{1}}\right\|.$
@@
$\left|\tilde{T}_{n_{1}}^{(P_{n},Q_{n})}-T_{n_{1}}^{(P_{n},Q_{n})}\right|\leq 2n_{1}\left\|\Sigma_{n_{1}}\right\|^{2}\left\{\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}+\left\|\Delta_{Y,Q_{n}}\right\|_{\mathrm{op}}+\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}\left\|\Delta_{Y,Q_{n}}\right\|_{\mathrm{op}}\right\}.$ (A.20)
We detail the $X$-part; the $Y$-part is identical with $(C_{YY},\mu_{j},\gamma^{Y}_{Q_{n}})$ in place of $(C_{XX},\lambda_{j},\gamma^{X}_{P_{n}})$.
@@
$A\coloneqq C_{XX},\qquad\hat{A}\coloneqq\hat{C}_{XX}=A+H,\qquad H\coloneqq\hat{C}_{XX}-C_{XX}.$
Let $\lambda_{1}>\lambda_{2}>\cdots$ be the (simple) positive eigenvalues of $A$ and fix $P_{n}$ such that
$\gamma^{X}_{P_{n}}\coloneqq\lambda_{P_{n}}-\lambda_{P_{n}+1}>0.$
Define the midpoint threshold
$r_{P_{n}}\coloneqq\frac{\lambda_{P_{n}}+\lambda_{P_{n}+1}}{2},\qquad\sigma_{X}\coloneqq(r_{P_{n}},\infty),\qquad\Sigma_{X}\coloneqq(-\infty,r_{P_{n}}].$
Note that
@@
$\mathrm{dist}\left(r_{P_{n}},\mathrm{spec}(A)\right)=\inf_{\lambda\in\mathrm{spec}(A)}\left|r_{P_{n}}-\lambda\right|=\min\left\{\lambda_{P_{n}}-r_{P_{n}},\,r_{P_{n}}-\lambda_{P_{n}+1}\right\}=\frac{\gamma^{X}_{P_{n}}}{2}.$
Let $E_{A}(\cdot)$ denote the spectral measure of the bounded self-adjoint operator $A$. Then the population projector onto the top-$P_{n}$ eigenspace can be written as the spectral projection
$P_{X,P_{n}}=E_{A}(\sigma_{X}).$
Define the “good” event
$\Omega^{X}_{n}\coloneqq\left\{\left\|H\right\|_{\mathrm{op}}<\frac{\gamma^{X}_{P_{n}}}{4}\right\}.$
On $\Omega^{X}_{n}$, Weyl’s eigenvalue inequality gives, for the ordered eigenvalues $\hat{\lambda}_{j}$ of $\hat{A}$,
$\hat{\lambda}_{P_{n}}\geq\lambda_{P_{n}}-\left\|H\right\|_{\mathrm{op}}>\lambda_{P_{n}}-\frac{\gamma^{X}_{P_{n}}}{4}=r_{P_{n}}+\frac{\gamma^{X}_{P_{n}}}{4}>r_{P_{n}},$
and
$\hat{\lambda}_{P_{n}+1}\leq\lambda_{P_{n}+1}+\left\|H\right\|_{\mathrm{op}}<\lambda_{P_{n}+1}+\frac{\gamma^{X}_{P_{n}}}{4}=r_{P_{n}}-\frac{\gamma^{X}_{P_{n}}}{4}<r_{P_{n}}.$
Hence, on $\Omega^{X}_{n}$, $\hat{A}$ has exactly $P_{n}$ eigenvalues in $(r_{P_{n}},\infty)$, so the empirical projector onto the top-$P_{n}$ eigenspace coincides with the spectral projection:
$\hat{P}_{X,P_{n}}=E_{\hat{A}}(\sigma_{X})\qquad\text{on }\Omega^{X}_{n}.$
Let $\Theta_{X}$ be the (operator) angle between the subspaces $\mathrm{Ran}(P_{X,P_{n}})$ and $\mathrm{Ran}(\hat{P}_{X,P_{n}})$. By the symmetric $\sin\Theta$ theorem of Davis and Kahan (1970),
$\mathrm{dist}\left(r_{P_{n}},\mathrm{spec}(A)\right)\,\left\|\sin\Theta_{X}\right\|_{\mathrm{op}}\leq\left\|H\right\|_{\mathrm{op}}.$
Moreover, for orthogonal projections one has the identity
$\left\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\right\|_{\mathrm{op}}=\left\|\sin\Theta_{X}\right\|_{\mathrm{op}}$
Therefore, on $\Omega^{X}_{n}$,
@@
$\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}=\left\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\right\|_{\mathrm{op}}\leq\frac{1}{\mathrm{dist}\left(r_{P_{n}},\mathrm{spec}(A)\right)}\left\|H\right\|_{\mathrm{op}}=\frac{2}{\gamma^{X}_{P_{n}}}\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{op}}.$ (A.21)
The same argument yields
$\left\|\Delta_{Y,Q_{n}}\right\|_{\mathrm{op}}\leq\frac{2}{\gamma^{Y}_{Q_{n}}}\left\|\hat{C}_{YY}-C_{YY}\right\|_{\mathrm{op}}.$ (A.22)
Since $\left\|\hat{P}_{X,P_{n}}-P_{X,P_{n}}\right\|_{\mathrm{op}}\leq 2$ always, for any $\varepsilon>0$,
@@
$\mathbb{P}\left(\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}>\varepsilon\right)\leq\mathbb{P}\left((\Omega^{X}_{n})^{c}\right)+\mathbb{P}\left(\frac{2}{\gamma^{X}_{P_{n}}}\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{op}}>\varepsilon\right).$
By Assumption 4.1, $\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{op}}/\gamma^{X}_{P_{n}}=o_{p}(1)$, hence both probabilities on the right-hand side converge to 0 and
$\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}=o_{p}(1).$
The same reasoning with (A.22) yields $\left\|\Delta_{Y,Q_{n}}\right\|_{\mathrm{op}}=o_{p}(1)$.
Next, we present a uniform convergence result. Fix $\varepsilon>0$. Using (A.21) and $\left\|\cdot\right\|_{\mathrm{op}}\leq\left\|\cdot\right\|_{\mathrm{HS}}$,
@@
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}>\varepsilon\right)\leq\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{HS}}>\frac{\varepsilon\,\gamma^{X}_{P_{n}}}{2}\right).$
By Markov’s inequality,
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{HS}}>\frac{\varepsilon\,\gamma^{X}_{P_{n}}}{2}\right)\leq\frac{2}{\varepsilon\,\gamma^{X}_{P_{n}}}\sup_{P\in\tilde{\mathcal{P}}_{0}}\operatorname{\mathbb{E}}_{P}\left[\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{HS}}\right].$
Under bounded kernels, the Hilbert–Schmidt norm of each centered rank-one summand in $\hat{C}_{XX}-C_{XX}$ is uniformly bounded, hence there exists a constant $C<\infty$ independent of $P\in\tilde{\mathcal{P}}_{0}$ such that
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\operatorname{\mathbb{E}}_{P}\left[\left\|\hat{C}_{XX}-C_{XX}\right\|_{\mathrm{HS}}\right]\leq\frac{C}{\sqrt{n_{1}}}.$
Therefore,
$\sup_{P\in\tilde{\mathcal{P}}_{0}}\mathbb{P}_{P}\left(\left\|\Delta_{X,P_{n}}\right\|_{\mathrm{op}}>\varepsilon\right)\leq\frac{2C}{\varepsilon\,\gamma^{X}_{P_{n}}\sqrt{n_{1}}}\longrightarrow 0$
by Assumption 4.5. The same argument applies to $\Delta_{Y,Q_{n}}$. Combining this with (A.20) and $\sqrt{n_{1}}\left\|\Sigma_{n_{1}}\right\|=O_{\tilde{\mathcal{P}}_{0}}(1)$ (which follows from boundedness of the kernels and the summands defining $\Sigma_{n_{1}}$), we obtain the claimed uniform convergence. ∎